# The Discovery of Little Red Dots in the Local Universe: Signatures of Cool Gas Envelopes

### v1

@@ D.3 The Haro 11 dust template
The inferred dust components are highly sensitive to the choice of fitting templates. In J1047+0739, the W2 photometry at a rest-frame wavelength of 4 µm is not well reproduced, as simple blackbody or graybody models fail to capture features such as PAH emission and silicate absorption. In J1022+0841, a dust component with K appears, but this primarily reflects the influence of the four dust templates adopted from J. Lyu & G. H. Rieke (2021). To explore alternative interpretations, we replace the dust component in Equation D1 with the Haro 11 dust template from J. Lyu et al. (2016) and fit the continua of the three local LRDs, as shown in Figure D.3. Notably, the dust emission in J1022+0841 is well reproduced by the Haro 11 template, suggesting that its dust may reside in the ISM and be heated by AGN-driven feedback or massive star clusters in the host galaxy. This degeneracy highlights the need for mid- to far-infrared observations to robustly constrain the dust content and its origin in these systems.

### v2

@@ D.3 The Haro 11 dust template
The inferred dust components are highly sensitive to the choice of fitting templates. In J1047+0739, the W2 photometry at a rest-frame wavelength of 4 µm is not well reproduced, as simple blackbody or graybody models fail to capture features such as PAH emission and silicate absorption. In J1022+0841, a dust component with K appears, but this primarily reflects the influence of the four dust templates adopted from Lyu2021. To explore alternative interpretations, we replace the dust component in Equation D1 with the Haro 11 dust template from Lyu2016 and fit the continua of the three local LRDs, as shown in Figure D.3. Notably, the dust emission in J1022+0841 is well reproduced by the Haro 11 template, suggesting that its dust may reside in the ISM and be heated by AGN-driven feedback or massive star clusters in the host galaxy. This degeneracy highlights the need for mid- to far-infrared observations to robustly constrain the dust content and its origin in these systems.